"Insert 3 rational numbers between -5 and -4" this is my problem pls solve it.
1
Expert's answer
2013-04-10T07:47:55-0400
You can insert infinitely many rational numbers between any two integers. For your task you pick, for example, the following
-5+1/2=-4.5=-9/5
-5+1/3=-4 2/3=-14/3
-5+1/4=-4 3/4=-19/4